# Equation Discovery with Bayesian Spike-and-Slab Priors and Efficient Kernels

### v1

@@ 4 Algorithm
Given the measurement data $\mathcal{D}$ and the dictionary ${\mathcal{O}}$, the learning of our model amounts to estimating ${\mathcal{U}}$, the kernel parameters, and the posterior distribution of ${\bf s}$ and ${\bf w}$. This is challenging. First, when the number of mesh points is large, the kernel interpolation that requires the inverse of ${\bf K}_{{\mathcal{M}},{\mathcal{M}}}$ can be extremely costly or even infeasible; see (5) and (6), because the time complexity is cubic to the number of mesh points. Second, because the spike-and-slab prior in (7) mixes binary and continuous random variables, the posterior distribution is analytically intractable. To address these challenges, we induce a Kronecker product in kernel computation, and we use tensor algebra to avoid operating on full matrices. We then develop an expectation-propagation expectation-maximization (EP-EM) algorithm. In each iteration, our algorithm performs two steps. In the E step, we fix ${\mathcal{U}}$ and the kernel parameters, and use EP (Minka, 2001a, ) to estimate the posterior of ${\bf s}$ and ${\bf w}$. In the M step, we maximize the expected model likelihood to update ${\mathcal{U}}$ and the kernel parameters. The alternating of the E and M steps keeps mutually improving equation discovery and function estimation until convergence. Our method is summarized in Algorithm 1.
@@
where ${\bf K}_{1}=\kappa(\bm{\gamma}_{1},\bm{\gamma}_{1})$, ${\bf K}_{2}=\kappa(\bm{\gamma}_{2},\bm{\gamma}_{2})$, and ${\bf K}_{3}=\kappa(\bm{\gamma}_{3},\bm{\gamma}_{3})$ are the kernel matrices of the inputs at each dimension of ${\mathcal{M}}$. Note that the popular square exponential (SE) kernel is a product kernel. One can also construct a product kernel from many other existing kernels.
@@
We can then leverage the properties of Kronecker product (Minka,, 2000) and tensor algebra (Kolda,, 2006) to avoid computing the full kernel matrix and tremendously improve efficiency. First, to compute (5), we can derive that
$u({\bf z})$ $=\kappa(t,\bm{\gamma}_{1})\otimes\kappa(x_{1},\bm{\gamma}_{2})\otimes\kappa(x_{2},\bm{\gamma}_{2})\left({\bf K}_{1}\otimes{\bf K}_{2}\otimes{\bf K}_{3}\right)^{-1}{\rm vec}({\mathcal{U}})$
$=\left(\kappa(t,\bm{\gamma}_{1}){\bf K}_{1}^{-1}\otimes\kappa(x_{1},\bm{\gamma}_{2}){\bf K}_{2}^{-1}\otimes\kappa({\bf x}_{2},\bm{\gamma}_{3}){\bf K}_{3}^{-1}\right){\rm vec}({\mathcal{U}})$
$={\mathcal{U}}\times_{1}\kappa(t,\bm{\gamma}_{1}){\bf K}_{1}^{-1}\times_{2}\kappa(x_{1},\bm{\gamma}_{2}){\bf K}_{2}^{-1}\times_{3}\kappa({\bf x}_{2},\bm{\gamma}_{3}){\bf K}_{3}^{-1},$ (11)
@@
where ×$\times_{k}$ is the tensor-matrix product at mode $k$ (Kolda,, 2006). Hence, we only need to compute the inverse of each ${\bf K}_{j}$ ($1\leq j\leq 3$), which takes time complexity $O(d_{1}^{3}+d_{2}^{3}+d_{3}^{3})$, and the tensor-matrix product at the last line of (11), which takes $O(d)$ time complexity ($d=d_{1}d_{2}d_{3}$) — linear in the size of ${\mathcal{U}}$. Hence, it is much more efficient than the naive computation — $O(d^{3})$ time complexity. Next, to compute a derivative at ${\mathcal{M}}$, we find that, because the product kernel is decomposed over individual input variables, the corresponding kernel differentiation still maintains the product structure. For example, to obtain ${\partial_{{x_{1}x_{2}}}}u$, we have the corresponding kernel derivative (see (6)) as
${\partial_{{x_{1}x_{2}}}}\kappa({\bf z},\bm{\rho}_{j})={\partial_{{x_{1}x_{2}}}}\left[\kappa(t,\rho_{j1})\kappa(x_{1},\rho_{j2})\kappa(x_{2},\rho_{j3})\right]=\kappa(t,\rho_{j1})\cdot{\partial_{{x_{1}}}}\kappa(x_{1},\rho_{j2})\cdot{\partial_{{x_{2}}}}\kappa(x_{2},\rho_{j3}).$
Accordingly, to compute ${\partial_{{x_{1}x_{2}}}}{\mathcal{U}}\overset{\Delta}{=}\{{\partial_{{x_{1}x_{2}}}}u({\bf z})|{\bf z}\in{\mathcal{M}}\}$, we can derive that
${\partial_{{x_{1}x_{2}}}}{\mathcal{U}}$ $=\left({\bf K}_{1}\otimes D_{1}[{\bf K}_{2}]\otimes D_{1}[{\bf K}_{3}]\right)\left({\bf K}_{1}\otimes{\bf K}_{2}\otimes{\bf K}_{3}\right)^{-1}{\rm vec}({\mathcal{U}})$
$=\left({\bf I}\otimes D_{1}[{\bf K}_{2}]{\bf K}_{2}^{-1}\otimes D_{1}[{\bf K}_{3}]{\bf K}_{3}^{-1}\right){\rm vec}({\mathcal{U}})$
$={\mathcal{U}}\times_{2}D_{1}[{\bf K}_{2}]{\bf K}_{2}^{-1}\times_{3}D_{1}[{\bf K}_{3}]{\bf K}_{3}^{-1},$ (12)
@@
where $D_{1}[\cdot]$ means taking the partial derivative w.r.t. the first input variable for each kernel function inside, $D_{1}[{\bf K}_{2}]=[{\partial_{{\gamma}}}\kappa(\gamma,\gamma^{\prime})]_{\gamma,\gamma^{\prime}\in\bm{\gamma}_{2}}$, and $D_{1}[{\bf K}_{3}]=[{\partial_{{\gamma}}}\kappa(\gamma,\gamma^{\prime})]_{\gamma,\gamma^{\prime}\in\bm{\gamma}_{3}}$. The multilinear operation has the time complexity $O((d_{2}+d_{3})d)$. Hence, the overall time complexity is $O(d_{1}^{3}+d_{2}^{3}+d_{3}^{3}+(d_{2}+d_{3})d)$ which again is much more efficient than the naive computation with $O(d^{3})$ complexity. Accordingly, we can compute the values of all the required derivatives on ${\mathcal{M}}$ highly efficiently.
@@
E Step. With this efficient computational method, we perform EP-EM steps for model estimation. In the E step, we estimate $p({\bf s},{\bf w}|\mathcal{D},{\widehat{{\bf y}}})$ given the current estimate of ${\mathcal{U}}$. According to (7) and (8), the joint distribution is
@@
$p({\bf w},{\bf s},\mathcal{D},{\widehat{{\bf y}}})\propto\prod\nolimits_{j=1}^{A}\text{Bern}(s_{j}|\rho_{0})\left({\bf s}_{j}\mathcal{N}(w_{j}|0,\sigma^{2}_{0})+(1-s_{m})\delta(w_{j})\right)\mathcal{N}({\bf h}|\bm{\Phi}{\bf w},\tau{\bf I}).$ (13)
@@
Note that we drop the constant terms to ${\bf w}$ and ${\bf s}$. We can see that the problem is reduced to the inference for Bayesian linear regression with the spike-and-slab prior. We therefore use a similar approach to (Fang et al.,, 2020) to develop an efficient EP method. EP approximates each non-exponential-family factor in the joint distribution with an exponential-family term. Then from the approximate joint distribution, we can obtain a closed-form posterior, because the exponential family is closed under multiplication. EP iteratively updates each factor approximation until convergence. We leave the details of our EP method in Section A of Appendix. After our EP inference, we obtain a posterior approximation,
@@
$p({\bf s},{\bf w}|\mathcal{D},{\widehat{{\bf y}}})\approx q({\bf s},{\bf w})=\prod\nolimits_{j=1}^{A}\text{Bern}(s_{j}|\sigma(\widehat{\rho}_{j}))\cdot\mathcal{N}({\bf w}|\bm{\beta},\bm{\Sigma}),$ (14)
@@
where $\sigma(\cdot)$ is the sigmoid function. We then prune the operators according to $q({\bf s})$ to decide the current equation.
@@
${\mathcal{L}}=-\frac{1}{2}{\rm vec}({\mathcal{U}})^{\top}{\bf K}_{{\mathcal{M}}{\mathcal{M}}}^{-1}{\rm vec}({\mathcal{U}})-\frac{1}{2v}\|{\bf y}-{\bf f}\|^{2}-\frac{1}{2\tau}\mathbb{E}_{q}\left[\left\|{\bf h}-\widehat{\bm{\Phi}}\widehat{{\bf w}}\right\|^{2}\right]+\text{const},$ (15)
where $\widehat{\bm{\Phi}}$ and $\widehat{{\bf w}}$ are the evaluation and weights of the remaining operators, respectively. Note that the first term is the RKHS norm of the target function estimate (5), which is used to regularize the learning of $u(\cdot)$. With the Kronecker product and tensor algebra, the computation of the RKHS norm is highly efficient. We can then use any gradient based optimization to maximize ${\mathcal{L}}$.
Algorithm Complexity. The time complexity of our model estimation algorithm is $O(\sum_{j=1}^{3}d_{j}^{3}+(N+\sum_{j}d_{j})d)$, where $d=d_{1}d_{2}d_{3}$. The space complexity is $O(\sum_{j=1}^{3}d_{j}^{2}+d+A^{2})$, which is to store the kernel matrices at each input dimension, function estimate ${\mathcal{U}}$ and $q({\bf s},{\bf w})$.
@@
3:     E-step: Fix ${\mathcal{U}}$ and $\bm{\theta}$, run EP to estimate $q({\bf s},{\bf w})$ in (14). Prune the operators with $q(s_{j}=1)<0.5$.
@@ 5 Related Work
Early work (Bongard and Lipson,, 2007; Schmidt and Lipson,, 2009) uses genetic programming and symbolic regression (Koza,, 1994) to find differential equations. These are costly and easily produce tediously long expressions. The recent break-through SINDy (Brunton et al.,, 2016) uses sparse linear regression to select PDE operators from a pre-specified dictionary. The sparsity is fulfilled by a sequential threshold ridge regression (STRidge) method, which repeatedly conducts least-mean-square estimation and weight truncation. While SINDy was originally developed to discover ODEs, by augmenting the dictionary with partial derivatives over spatial variables, SINDy can be extended for discovering PDEs for spatial-temporal systems (Rudy et al.,, 2017; Schaeffer,, 2017). The SINDy family of approaches use numerical differentiation to evaluate candidate operators, and it can suffer from scarce and noisy data. To alleviate this issue, recent works (Berg and Nyström,, 2019; Both et al.,, 2021; Chen et al., 2021b, ) use deep NNs to approximate the solution, applying automatic differentiation to evaluate and select the operators. These methods can therefore be viewed as instances/extension of the PINNs (Raissi et al.,, 2019; Krishnapriyan et al.,, 2021). In (Berg and Nyström,, 2019), a deep NN is first used to fit the data, and then a sparse linear regression with $L_{1}$ regularization is applied to select the operators. In (Both et al.,, 2021), the NN and sparse linear regression are jointly trained. In (Chen et al., 2021b, ), a joint training strategy is also used, but it performs an alternating optimization of the NN loss and sparse regression, and it uses STRidge for weight truncation. The most recent work (Sun et al.,, 2022) develops a Bayesian spline learning (BSL) method, which uses splines to approximate the solution and a student-$t$ prior, i.e., relevance vector machines (Tipping,, 2001), for sparse linear regression. It uses a similar alternating optimization strategy to (Chen et al., 2021b, ). To quantify the uncertainty, BSL then applies Stochastic Weight Averaging-Gaussian (SWAG) (Maddox et al.,, 2019) for posterior approximation of the weights.
@@
Our work uses kernel methods to estimate the solution function for PDE discovery. GP/kernel methods have been applied for solving differential equations for a long time. For example, Graepel, (2003); Raissi et al., (2017) used GPs to solve linear PDEs with the noisy source term measurements. The recent work (Chen et al., 2021a, ) developed a general kernel method to solve linear/nonlinear PDEs. However, all these methods assume the equations are given, and thus they cannot discover the equations from data. GPs have also been used to model the equation components or incorporate the equations for better training, such as (Heinonen et al.,, 2018; Alvarez et al.,, 2009; Barber and Wang,, 2014; Macdonald et al.,, 2015; Heinonen et al.,, 2018; Wenk et al.,, 2019, 2020; Long et al.,, 2022). The GP community has realized the computational advantage of the Kronecker product (Saatcci,, 2012), and there has been work in leveraging the Kronecker product properties to improve the training speed and scalability, such as high-dimension output regression (Zhe et al.,, 2019) and sparse approximation based on inducing points (Wilson and Nickisch,, 2015) However, in typical machine learning applications, the data are not observed at a grid and the Kronecker product structure has a limited usage. By contrast, for PDE solving, it is natural to estimate the solution values on a mesh (which is consistent with the practice of traditional numerical methods). This opens the possibility of using Kronecker products for efficient computation. To our knowledge, our work is first to realize this point and use the Kronecker product structure to enable efficient PDE discovery and solution estimation.

### v2

@@ 4 ALGORITHM
Given the measurement data $\mathcal{D}$ and the dictionary ${\mathcal{O}}$, the learning of our model amounts to estimating ${\mathcal{U}}$, the kernel parameters, and the posterior distribution of ${\bf s}$ and ${\bf w}$. This is challenging. First, when the number of mesh points is large, the kernel interpolation that requires the inverse of ${\bf K}_{{\mathcal{M}},{\mathcal{M}}}$ can be extremely costly or even infeasible; see (4) and (5). Second, because the spike-and-slab prior (2) mixes binary and continuous random variables, the posterior distribution is analytically intractable. To address these challenges, we induce a Kronecker product in kernel computation, and use tensor algebra to avoid operating on full matrices. We then develop an expectation-propagation expectation-maximization (EP-EM) algorithm. In each iteration, our algorithm performs two steps. In the E step, we fix ${\mathcal{U}}$ and the kernel parameters, and use EP (Minka, 2001a, ) to estimate the posterior of ${\bf s}$ and ${\bf w}$. In the M step, we maximize the expected model likelihood to update ${\mathcal{U}}$ and the kernel parameters. The alternating of the E and M steps keeps mutually improving equation discovery and function estimation until convergence. See Algorithm 1 for a summary.
@@
where ${\bf K}_{1}=\kappa(\bm{\gamma}_{1},\bm{\gamma}_{1})$, ${\bf K}_{2}=\kappa(\bm{\gamma}_{2},\bm{\gamma}_{2})$, and ${\bf K}_{3}=\kappa(\bm{\gamma}_{3},\bm{\gamma}_{3})$ are the kernel matrices of the inputs at each dimension of ${\mathcal{M}}$. Note that the popular square exponential (SE) kernel is a product kernel. One can also construct a product kernel from any other kernels.
@@
We can then leverage the nice properties of Kronecker product (Minka,, 2000) and tensor algebra (Kolda,, 2006) to avoid computing the full kernel matrix and tremendously improve the efficiency. First, to compute (4), we can derive that $u({\bf z})=\kappa(t,\bm{\gamma}_{1})\otimes\kappa(x_{1},\bm{\gamma}_{2})\otimes\kappa(x_{2},\bm{\gamma}_{3})\left({\bf K}_{1}\otimes{\bf K}_{2}\otimes{\bf K}_{3}\right)^{-1}{\rm vec}({\mathcal{U}})=\left(\kappa(t,\bm{\gamma}_{1}){\bf K}_{1}^{-1}\otimes\kappa(x_{1},\bm{\gamma}_{2}){\bf K}_{2}^{-1}\otimes\kappa(x_{2},\bm{\gamma}_{3}){\bf K}_{3}^{-1}\right){\rm vec}({\mathcal{U}})={\mathcal{U}}\times_{1}\kappa(t,\bm{\gamma}_{1}){\bf K}_{1}^{-1}\times_{2}\kappa(x_{1},\bm{\gamma}_{2}){\bf K}_{2}^{-1}\times_{3}\kappa(x_{2},\bm{\gamma}_{3}){\bf K}_{3}^{-1}$, where ×$\times_{k}$ is the tensor-matrix product at mode $k$ (Kolda,, 2006). Hence, we only need to compute the inverse of each ${\bf K}_{j}$ ($1\leq j\leq 3$), which takes time complexity $O(d_{1}^{3}+d_{2}^{3}+d_{3}^{3})$, and the tensor-matrix product takes $O(d)$ time complexity ($d=d_{1}d_{2}d_{3}$) — linear in the size of ${\mathcal{U}}$. Hence, it is much more efficient than the naive computation — $O(d^{3})$ time complexity.
@@
Next, to compute a derivative of $u$ at ${\mathcal{M}}$, we find that, because the product kernel is decomposed over individual input variables, the corresponding kernel differentiation still maintains the product structure. For example, to obtain ${\partial_{{x_{1}x_{2}}}}u$, we have the corresponding kernel derivative (see (5)) as
${\partial_{{x_{1}x_{2}}}}\kappa({\bf z},\bm{\rho}_{j})={\partial_{{x_{1}x_{2}}}}\left[\kappa(t,\rho_{j1})\kappa(x_{1},\rho_{j2})\kappa(x_{2},\rho_{j3})\right]$
$=\kappa(t,\rho_{j1})\cdot{\partial_{{x_{1}}}}\kappa(x_{1},\rho_{j2})\cdot{\partial_{{x_{2}}}}\kappa(x_{2},\rho_{j3}).$ (9)
Accordingly, to compute ${\partial_{{x_{1}x_{2}}}}{\mathcal{U}}\overset{\Delta}{=}\{{\partial_{{x_{1}x_{2}}}}u({\bf z})|{\bf z}\in{\mathcal{M}}\}$, we have
${\rm vec}({\partial_{{x_{1}x_{2}}}}{\mathcal{U}})$
$=\left({\bf K}_{1}\otimes D_{1}[{\bf K}_{2}]\otimes D_{1}[{\bf K}_{3}]\right)\left({\bf K}_{1}\otimes{\bf K}_{2}\otimes{\bf K}_{3}\right)^{-1}{\rm vec}({\mathcal{U}})$
$=\left({\bf I}\otimes D_{1}[{\bf K}_{2}]{\bf K}_{2}^{-1}\otimes D_{1}[{\bf K}_{3}]{\bf K}_{3}^{-1}\right){\rm vec}({\mathcal{U}}),$
and hence
${\partial_{{x_{1}x_{2}}}}{\mathcal{U}}={\mathcal{U}}\times_{2}D_{1}[{\bf K}_{2}]{\bf K}_{2}^{-1}\times_{3}D_{1}[{\bf K}_{3}]{\bf K}_{3}^{-1},$ (10)
@@
where $D_{1}[\cdot]$ means taking the partial derivative w.r.t the first input variable for each kernel function inside, $D_{1}[{\bf K}_{2}]=[{\partial_{{\gamma}}}\kappa(\gamma,\gamma^{\prime})]_{\gamma,\gamma^{\prime}\in\bm{\gamma}_{2}}$, and $D_{1}[{\bf K}_{3}]=[{\partial_{{\gamma}}}\kappa(\gamma,\gamma^{\prime})]_{\gamma,\gamma^{\prime}\in\bm{\gamma}_{3}}$. The multilinear operation takes time complexity $O((d_{2}+d_{3})d)$. Hence, the overall time complexity is $O(d_{1}^{3}+d_{2}^{3}+d_{3}^{3}+(d_{2}+d_{3})d)$ which again is much more efficient than the naive computation with $O(d^{3})$ complexity. As such, we can compute the values of all the required derivatives on ${\mathcal{M}}$ highly efficiently.
@@
E Step. With this efficient computational method, we perform EP-EM steps for model estimation. In the E step, we estimate $p({\bf s},{\bf w}|\mathcal{D},{\widehat{{\bf y}}})$ given the current estimate of ${\mathcal{U}}$. The joint distribution is
$p({\bf w},{\bf s},\mathcal{D},{\widehat{{\bf y}}})\propto\mathcal{N}({\bf h}|\bm{\Phi}{\bf w},\tau{\bf I})\cdot$ (11)
@@
$\hskip 2.84526pt\prod\nolimits_{j=1}^{A}\text{Bern}(s_{j}|\rho_{0})\left({\bf s}_{j}\mathcal{N}(w_{j}|0,\sigma^{2}_{0})+(1-s_{m})\delta(w_{j})\right).$
@@
Note that we drop the constant terms to ${\bf w}$ and ${\bf s}$, such as the measurement data likelihood (7). We can see that the problem is reduced to the inference for Bayesian linear regression with the spike-and-slab prior. We therefore use a similar approach to (Fang et al.,, 2020) to develop an efficient EP method. EP approximates each non-exponential-family factor in the joint distribution with an exponential-family term. Then from the approximate joint distribution, we can obtain a closed-form posterior, since the exponential family is closed under multiplication. EP iteratively updates each factor approximation until convergence. We leave the details in Appendix Section A.1. After our EP inference, we obtain a posterior approximation,
$p({\bf s},{\bf w}|\mathcal{D},{\widehat{{\bf y}}})$
@@
$\approx q({\bf s},{\bf w})=\prod\nolimits_{j=1}^{A}\text{Bern}(s_{j}|\sigma(\widehat{\rho}_{j}))\cdot\mathcal{N}({\bf w}|\bm{\beta},\bm{\Sigma}),$ (12)
@@
where $\sigma(\cdot)$ is the sigmoid function. We then prune the operators according to $q({\bf s})$ to decide the currently discovered equation.
@@
${\mathcal{L}}=$ $-\frac{1}{2}{\rm vec}({\mathcal{U}})^{\top}{\bf K}_{{\mathcal{M}}{\mathcal{M}}}^{-1}{\rm vec}({\mathcal{U}})-\frac{1}{2v}\|{\bf y}-{\bf f}\|^{2}$
$-\frac{1}{2\tau}\mathbb{E}_{q}\left[\left|{\bf h}-\widehat{\bm{\Phi}}\widehat{{\bf w}}\right\|^{2}\right]+\text{const},$ (13)
where $\widehat{\bm{\Phi}}$ and $\widehat{{\bf w}}$ are the evaluation and weights of the remaining operators, respectively. Note that the first term is the RKHS norm of the target function estimate (4), which regularizes the learning of $u(\cdot)$. With the Kronecker product and tensor algebra, the computation of the RKHS norm is highly efficient. We can use any gradient based optimization to maximize ${\mathcal{L}}$.
Algorithm Complexity. The time complexity of our model estimation algorithm is $O(\sum_{j=1}^{3}d_{j}^{3}+(N+\sum_{j}d_{j})d)$, where $d=d_{1}d_{2}d_{3}$. The space complexity is $O(\sum_{j=1}^{3}d_{j}^{2}+d+A^{2})$, for the storage of the kernel matrices at each dimension, ${\mathcal{U}}$ and $q({\bf s},{\bf w})$.
@@
3:     E-step: Fix ${\mathcal{U}}$ and $\bm{\theta}$, run EP to estimate $q({\bf s},{\bf w})$ in (12). Prune the operators with $q(s_{j}=1)<\alpha$.
@@ 5 RELATED WORK
The recent break-through SINDy (Brunton et al.,, 2016) uses sparse linear regression to select PDE operators from a pre-specified dictionary. The sparsity is fulfilled by a sequential threshold ridge regression (STRidge) method, which repeatedly conducts least-mean-square estimation and weight truncation. While SINDy was originally developed to discover ODEs, by augmenting the dictionary with partial derivatives over spatial variables, SINDy can be extended for discovering PDEs for spatial-temporal systems (Rudy et al.,, 2017; Schaeffer,, 2017). The SINDy family of approaches use numerical differentiation to evaluate candidate operators, and it can suffer from scarce and noisy data. To alleviate this issue, recent works (Berg and Nyström,, 2019; Both et al.,, 2021; Chen et al., 2021b, ) use deep NNs to approximate the solution, applying automatic differentiation to evaluate and select the operators. These methods can therefore be viewed as instances or extension of PINNs (Raissi et al.,, 2019; Krishnapriyan et al.,, 2021). In (Berg and Nyström,, 2019), a deep NN is first used to fit the data, and then a sparse linear regression with $L_{1}$ regularization is applied to select the operators. In (Both et al.,, 2021), the NN and sparse linear regression are jointly trained. In (Chen et al., 2021b, ), a joint training strategy is also used, but it performs an alternating optimization of the NN loss and sparse regression, and it uses STRidge for weight truncation. The most recent work (Sun et al.,, 2022) develops a Bayesian spline learning (BSL) method, which uses splines to approximate the solution and a student-$t$ prior, i.e., relevance vector machines (Tipping,, 2001), for sparse linear regression. It uses a similar alternating optimization strategy to (Chen et al., 2021b, ). To quantify the uncertainty, BSL then applies Stochastic Weight Averaging-Gaussian (SWAG) (Maddox et al.,, 2019) for posterior approximation of the weights.
@@
Our work uses kernel methods to estimate the solution function for PDE discovery. GP or kernel methods have been applied for solving differential equations for a long time. For example, Graepel, (2003); Raissi et al., (2017) used GPs to solve linear PDEs with the noisy source term measurements. The recent work (Chen et al., 2021a, ) developed a general kernel method to solve nonlinear PDEs. Fang et al., (2023) developed a spectral mixture kernel for solving high frequency and multi-scale PDEs. However, all these methods assume the equations are given, and thus they cannot discover the equations from data. GPs have also been used to model the equation components or incorporate the equations for better training, such as (Heinonen et al.,, 2018; Alvarez et al.,, 2009; Barber and Wang,, 2014; Macdonald et al.,, 2015; Wenk et al.,, 2019, 2020). Long et al., (2022) proposed a general framework to integrate PDEs/ODEs into GP training. The GP community has realized the computational advantage of the Kronecker product (Saatcci,, 2012), and there has been work in leveraging the Kronecker product properties to improve the training speed and scalability, such as high-dimension output regression (Zhe et al.,, 2019; Li et al.,, 2021; Wang et al.,, 2021) and sparse approximation based on inducing points (Wilson and Nickisch,, 2015). However, in typical machine learning applications, the data are not observed at a grid and the Kronecker product has a limited usage. By contrast, for PDE solving, it is natural to estimate the solution values on a mesh (which is consistent with the practice of traditional numerical methods). This opens the possibility of using Kronecker products to enable efficient PDE solution estimation.